Calculate the time period of an object having mass moment of inertia = 100 Kg-m^2, mass of 10 Kg and the centre of gravity lies at a point 20 cm below the point of suspension.
1 : 14.1 s
2 : 15.2 s
3 : 13.3 s
4 : 12.9 s
1
Expert's answer
2020-08-02T15:07:42-0400
"T=2\pi\sqrt{\frac{I}{mgL_{COM}}}"

Thus,

"T=2\pi\sqrt{\frac{100}{10\times 9.8\times0.2}}=14.1s"

Correct option is (1).